3. How many grams of aluminum were used to prepare 20.0 g of potassium alum if the percent
yield was 81.5%?
2 KAl(OH)4 + 4 H2SO4 → K2SO4 + Al2(SO4)3 + 8 H2O
1
Expert's answer
2020-11-26T06:02:52-0500

The reaction to get the potassium alum from aluminium:

"2Al + 2KOH + 22 H_2O + 4H_2SO_4= 2KAl(SO_4)_2 + 12H_2O+3H_2"


The number of mols of Al equals to number of mols of "KAl(SO4)_2" :

"n(Al) = n(KAl(SO_4)_2)"

"n(KAl(SO_4)_2) = \frac{m(KAl(SO_4)_2)}{M(KAl(SO_4)_2)}= \frac{20.0 g}{258.2 g/mol}=0.077 mol"


From this, the mass of aluminium equals to:

"m(Al) = n(Al)\cdot M(Al) = 0.077 mol \cdot 28 g/mol = 2.17 g"


By taking into account a knowledge about percentage yield, we can compute the real mass of aluminium reacted to form potassium alum:

"m(Al) = 2.17 g \cdot 0.815 = 1.77 g"